# ASTRID - An Automated and Scalable TRIaD for the Evaluation of RAG-based Clinical Question Answering Systems

### v1

@@ Abstract
Large Language Models (LLMs) have shown impressive potential in clinical question answering (QA), with Retrieval Augmented Generation (RAG) emerging as a leading approach for ensuring the factual accuracy of model responses. However, current automated RAG metrics perform poorly in clinical and conversational use cases. Using clinical human evaluations of responses is expensive, unscalable, and not conducive to the continuous iterative development of RAG systems. To address these challenges, we introduce ASTRID - an Automated and Scalable TRIaD for evaluating clinical QA systems leveraging RAG - consisting of three metrics: Context Relevance (CR), Refusal Accuracy (RA), and Conversational Faithfulness (CF). Our novel evaluation metric, CF, is designed to better capture the faithfulness of a model’s response to the knowledge base without penalising conversational elements. To validate our triad, we curate a dataset of over 200 real-world patient questions posed to an LLM-based QA agent during surgical follow-up for cataract surgery - the highest volume operation in the world - augmented with clinician-selected questions for emergency, clinical, and non-clinical out-of-domain scenarios. We demonstrate that CF can predict human ratings of faithfulness better than existing definitions for conversational use cases. Furthermore, we show that evaluation using our triad consisting of CF, RA, and CR exhibits alignment with clinician assessment for inappropriate, harmful, or unhelpful responses. Finally, using nine different LLMs, we demonstrate that the three metrics can closely agree with human evaluations, highlighting the potential of these metrics for use in LLM-driven automated evaluation pipelines. We also publish the prompts and datasets for these experiments, providing valuable resources for further research and development.
@@ 1 Introduction
The healthcare industry is increasingly adopting automation to meet rising demands on resources [33]. Large Language Models (LLMs) due to their capabilities have become increasingly popular in supportive clinical applications such as note-taking and summarisation[3]. A crucial aspect of patient care is the ability to ask questions and receive answers, which has been enhanced by advancements in Question-Answering (QA) systems powered by LLMs. However, the issue of hallucination remains a significant barrier in using LLMs for clinical QA systems [32]. Retrieval Augmented Generation (RAG) is a technique developed to address hallucination and ensure context appropriateness [21]. Despite these advancements, RAG systems lack sufficient evaluation metrics and frameworks, making it difficult to quantitatively establish their safety and identify system deficiencies.
This work explores the limitations of current evaluation methods and applies safety engineering principles to identify potential hazard cases in clinical QA [12, 8]. We develop a robust and scalable framework of metrics to systematically demonstrate how developers can mitigate potential hazards in LLM-based QA systems for clinical use. Using real patient questions from clinical trials on cataract post-operative recovery, we illustrate how these metrics can be interpreted in a clinical context. We validate our metrics by proving they model human ratings better than previous metrics, and effectively predict clinical harm, usefulness, and inappropriateness as labelled by specialist doctors. Our aim is to establish a foundation for developing and assessing LLM-powered clinical QA systems and encourage further research in this area.
@@
Our contributions are summarised as follows (Figure 4):
  • A hazard analysis of clinical QA systems inspired by the safety engineering principles.
@@
  • An analysis of these metrics and how they model human ratings.
  • An analysis of how these metrics can predict clinical harm, usefulness, and inappropriateness to a high standard when used together.
Figure 2: ASTRID - an Automated and Scalable TRIaD for evaluating clinical QA systems leveraging RAG - consisting of three metrics: Context Relevance (CR), Refusal Accuracy (RA), and Conversational Faithfulness (CF) assessed within a clinical context.
@@ 2.1 Background to clinical QA evaluation
Clinical QA systems powered by LLMs have generated significant recent interest. Already, some LLMs have demonstrated capabilities to generate more accurate responses [42, 2, 36, 48, 46], and sometimes even more empathetic [20] than doctors across various clinical contexts. However, LLMs can generate plausible-sounding, but factually incorrect responses, commonly referred to as ‘hallucinations’ [15]. Moreover, LLMs have a cut-off date when it comes to their knowledge [29] and this can pose significant safety risks in healthcare. While these issues can be somewhat addressed using RAG, demonstrating they are addressed is still a challenge.
To evaluate some of these risks specific to clinical QA systems using RAG, various efforts have been made to develop performance benchmarks. Currently, published benchmarks often utilise multiple-choice or categorical ground-truth answers for responses [49, 22, 47, 28], which fail to capture the complexities and risks associated with open-ended response generations. Where open-ended answers are evaluated, n-gram-based metrics such as BLEU [30], or ROUGE [23], historically used for machine translation, have been used [4]. However, these evaluations have been criticised for failing to capture the nuanced requirements of clinical QA, and even transformer-based metrics such as BertScore [52] have numerous semantic limitations [6].
A key feature of these risks in the context of open-ended clinical QA is their non-bimodal nature (i.e. an answer is not "safe" or not on a single axis). Consequently, the gold standard for assessing clinical inappropriateness remains human evaluation. For instance, Google’s work in clinical QA involved both clinicians and lay individuals to label responses on various axes such as the likelihood and severity of harm, alignment with scientific consensus, and helpfulness [37]. Similarly, other studies have employed multi-axis evaluations with human clinicians to assess the overall appropriateness of responses for open-ended clinical QA [26, 38, 51, 5].
@@ 2.2 Current RAG metrics
Evaluating RAG systems presents challenges due to their hybrid structure and the overall quality of the output often depends on multiple components within the systems. While attempts have been made to assess the overall quality of responses using deterministic methods [24, 25], most of the current evaluation metrics for RAG systems use an ensemble of component-level assessments, the majority of which leverage LLMs as judges [50]. Broadly, the performance of RAG pipelines can be evaluated by examining two main components: the retrieval and the generation components. For the retrieval component, key metrics include context relevance and retrieval accuracy. For the generation component, such metrics include answer relevance, faithfulness, and answer correctness.
These component evaluations have been variably implemented with popular tools including TruEra’s RAG Triad [45], and LangChain Bench [19]. Additionally, LLM-as-a-judge-based frameworks like RAGAS [9], and ARES [34] have popularised common evaluation triads to capture possible permutations of the above components. Please see appendix A.1 for an example on how the three components of the RAG system can be judged by LLMs, using the RAGAS metrics as an example.
@@ 2.2.1 Limitations of current metrics
The established methods to measure Faithfulness break down a model’s response into granular statements and then evaluate each statement’s consistency with the context [9]. This approach aims to create more focused assertions that consider the context of both the question and the answer. It is particularly advantageous when answers are short and lack context when reviewed in isolation, as demonstrated by Figure 6. However, in the context of clinical conversations, this approach has some shortcomings. Firstly, summarising responses into statements often overlooks the clinical nuances present in the original dialogue (Figure 7). Creating statements from both the patient’s question and the agent’s answer can hinder an independent review of the agent’s response in relation to the context. This is particularly problematic when the combined statements contain factual inaccuracies (Figure 8). Lastly, dialogue agents, particularly in clinical settings, are prompted to respond pathetically and conversationally. Statements constructed from the agent’s acknowledgments and questions, such as those meant to clarify or follow up on the patient’s queries or concerns, are penalised by existing faithfulness definitions (Figure 9).
@@
Evaluating answer relevance is critical in QA systems to ensure generated responses align with query intent. However, most current definitions focus on lexical or semantic similarity between the question and the response [9, 39]. Such approaches over-emphasise surface-level topic matching without accounting for deeper contextual understanding. Additionally, they neglect to factor in whether a context is appropriate given a clinical context.
In a conversational context, a simple answer such as "yes" or "no" could be entirely appropriate, and constitute a clinically meaningful (and thus risky) response, which will not be captured by answer relevance metrics.
Additional Limitations
Furthermore, existing metrics often penalise the system for appropriately refusing to address a question when it falls outside its scope of relevance or when there is insufficient information to provide a safe and accurate response. This is crucial as clinical QA systems are often required to stay within the defined scope of practice.
@@ 3.1 Deriving metrics towards a safety case
In order to align our framework towards the evidence required to demonstrate if a clinical system is safe, we sought inspiration from published safety engineering frameworks - namely the Safety Assurance of autonomous systems in Complex Environments (SACE) guidance [12]. Structured safety engineering approaches have been applied towards the assurance of high-integrity autonomous systems (AS) such as maritime vessels [27], automotive [31, 14], aerospace [43], and healthcare domains [16, 10].
@@
The SACE framework, in particular, provides a process to systematically integrate safety assurance into the development of AS whilst considering the system and its environment. Whilst we do not report all artefacts from the process in its entirety, we highlight a few key steps in this process that have been applied towards ASTRID’s design. Namely, we considered the principles of:
  • Operating Context Assurance: What are the different clinical scenarios that a patient could conceivably pose to a clinical QA agent? (Figure 4)
  • Hazardous Scenario Identification: How can RAG systems behave in hazardous ways in each of these scenarios?
  • Safe Operating Concept Assurance: How should an ideal system behave?
  • Out of Context Operation Assurance: What should a safe response be when a question is asked out of the clinical context for that interaction?
Figure 4: Clinical Operating Contexts that face a clinical QA agent.
We observe that the environment (clinical context) is crucial to answer safety. For example, for the question "Is it normal to have stomach cramps and vomiting?", this question in the context of a follow up appointment for routine eye surgery is unlikely to be relevant, and one would expect the system to not respond. However, if this was in the context of a patient who has just gone home following bowel surgery, this is likely to not only be highly relevant, but one would expect the system to respond (Figure 3).
These concepts were outlined in a workshop where the dataset of real-world questions posed by patients to a voice-based conversational Artificial Intelligence (AI) were reviewed. The workshop consisted of two AI developers, a clinician and safety practitioner (summarised in Appendix A.4), and the analysis provided a bridge between subjective clinical assessments of harms and helpfulness, with component-level validation scenarios for appropriate system performance.
@@ 3.2 A novel set of metrics and a framework to assess safety risks
Current RAG metrics do not correlate to clinical risks, and have varying levels of validation against human evaluations, with poor performance in conversational contexts. To our knowledge, there have also been no efforts to connect QA system performance with automated metrics for RAG systems, with real-world clinician grading of clinical harms, helpfulness and inappropriateness of responses. For developers to meaningfully understand whether a clinical RAG QA system meets safe operating concepts, we needed a framework that was validated for clinical use, scalability, and acknowledged nuanced clinical contexts.
We propose a novel Automated and Scalable Triad (ASTRID) analysis framework for RAG-based clinical QA systems. ASTRID consists of three reference-free LLM-based metrics: Refusal Accuracy (RA), Conversational Faithfulness (CF) and Context Relevance (CR) (Figure 2). In the subsequent sections, we will illustrate how to validate each of the metrics and the overall framework based on real-world data from patients speaking to clinical conversational agents, augmented to ensure sufficient test-case coverage.
@@ 3.2.1 Conversational Faithfulness (CF)
Evaluating how grounded a response is concerning the information provided is important to QA systems using RAG. Existing metrics that address this do not encapsulate additional complexities associated with conversational agents in a clinical setting. Therefore we propose a new metric Conversational Faithfulness (CF).
Given an answer-context pair, Conversational Faithfulness is defined as the proportion of information-containing sentences that are faithful to the context. To calculate CF, we employ the following steps:
@@
    We categorise different sentences in the response into "acknowledgements", "questions" and "informative". We provide the prompt used to achieve this step in the appendix (1).
@@
    We determine whether the informative sentences are grounded in context. We provide the prompt used to achieve this step in the appendix (2).
Finally, CF is calculated as the fraction of the number of informative sentences grounded in context to the total number of informative sentences.
@@ 3.2.2 Refusal Accuracy (RA)
As discussed in previous sections, an important aspect of evaluating QA systems in the clinical setting is the ability of the system to decline to respond when it cannot answer a question, or a question is not appropriate for the clinical context. This is essential especially in LLM-powered systems, where risks arise from a model’s tendency to provide ungrounded responses. As current metrics do not capture this behaviour, we add the metric Refusal Accuracy (RA) to our triad.
Refusal Accuracy is defined as the system’s ability to deny a response when there is no relevant information available to answer the question. We use binary labels to indicate whether the system appropriately refuses to respond. We provide the prompt used to achieve this step in the appendix (3).

### v2

@@ Abstract
Large Language Models (LLMs) have shown impressive potential in clinical question answering (QA), with Retrieval Augmented Generation (RAG) emerging as a leading approach for ensuring the factual accuracy of model responses. However, current automated RAG metrics perform poorly in clinical and conversational use cases. Using clinical human evaluations of responses is expensive, unscalable, and not conducive to the continuous iterative development of RAG systems. To address these challenges, we introduce ASTRID - an Automated and Scalable TRIaD for evaluating clinical QA systems leveraging RAG - consisting of three metrics: Context Relevance (CR), Refusal Accuracy (RA), and Conversational Faithfulness (CF). Our novel evaluation metric, CF, is designed to better capture the faithfulness of a model’s response to the knowledge base without penalizing conversational elements. Additionally, our metric RA captures the refusal to address questions outside of the system’s scope of practice. To validate our triad, we curate a dataset of over 200 real-world patient questions posed to an LLM-based QA agent during surgical follow-up for cataract surgery - the highest volume operation in the world - augmented with clinician-selected questions for emergency, and clinical and non-clinical out-of-domain scenarios. We demonstrate that CF predicts human ratings of faithfulness more accurately than existing definitions in conversational settings. Furthermore, using eight different LLMs, we demonstrate that the three metrics can closely agree with human evaluations, highlighting the potential of these metrics for use in LLM-driven automated evaluation pipelines. Finally, we show that evaluation using our triad of CF, RA, and CR exhibits alignment with clinician assessment for inappropriate, harmful, or unhelpful responses. We also publish the prompts and datasets for these experiments, providing valuable resources for further research and development.
ASTRID - An Automated and Scalable TRIaD for the Evaluation of RAG-based Clinical Question Answering Systems
Yajie Vera He [^2] Authors contributed equally. Ufonia Limited yh@ufonia.com                        Mohita Chowdhury [^2] Authors contributed equally. [^1] Now at Google DeepMind. Ufonia Limited chowdhury.mohita@gmail.com                        Jared Joselowitz Ufonia Limited jj@ufonia.com
Aisling Higham Ufonia Limited ah@ufonia.com                        Ernest Lim Ufonia Limited / University of York el@ufonia.com
@@ 1 Introduction
The healthcare industry is increasingly adopting automation to meet rising demands on resources Ruiz and Duffy (2021). LLMs, due to their capabilities, have become increasingly popular in supportive clinical applications such as note-taking and summarization Cascella et al. (2023). A crucial aspect of patient care is the ability to ask questions and receive answers, which has been enhanced by advancements in QA systems powered by LLMs. However, the issue of hallucination remains a significant barrier to using LLMs for clinical QA systems Rawte et al. (2023). RAG is a technique developed to address hallucination and ensure context appropriateness Lewis et al. (2020). Despite these advancements, RAG systems lack sufficient evaluation metrics and frameworks, making it difficult to quantitatively establish their safety and identify system deficiencies. Figure 1 illustrates the limitations of current clinical evaluation approaches and how automated methods address these challenges. This work examines evaluation limitations and applies safety engineering to identify hazard cases in clinical QA Hawkins et al. (2022); Ericson et al. (2015). We develop a robust, scalable framework of metrics to systematically demonstrate how developers can mitigate potential hazards in LLM-based QA systems for clinical use. Using real patient questions from clinical trials on cataract post-operative recovery, we illustrate how these metrics can be interpreted in a clinical context. We validate our metrics by proving they model human ratings better than previous metrics and effectively predict clinical harm, usefulness, and inappropriateness as labeled by specialist doctors. Our aim is to establish a foundation for developing and assessing LLM-powered clinical QA systems and encourage further research in this area.
@@
Our contributions are summarized as follows:
  • A hazard analysis of clinical QA systems inspired by safety engineering principles.
@@
  • A formal evaluation of these metrics and their alignment with human ratings.
  • An analysis of how these metrics collectively predict clinical harm, usefulness, and inappropriateness with high accuracy.
  • An assessment of how these metrics can be automated across eight different LLMs.
@@ 2.1 Background to clinical QA evaluation
Clinical QA systems powered by LLMs have generated significant recent interest. Already, some LLMs have demonstrated capabilities to generate more accurate responses Thirunavukarasu et al. (2023); Bernstein et al. (2023); Samaan et al. (2023); Xie et al. (2023); Van Bulck and Moons (2024), and sometimes even more empathetic than doctors across various clinical contexts Lee et al. (2024). However, LLMs can generate plausible-sounding, but factually incorrect responses, commonly referred to as "hallucinations" Ji et al. (2023). Moreover, LLMs have knowledge cut-off date Ovadia et al. (2023) and this can pose significant safety risks in healthcare. While these issues can be mitigated using RAG, risks still remain.
To address some of these risks specific to clinical QA systems using RAG, various efforts have been made to develop performance benchmarks. Currently, published benchmarks often utilize multiple-choice or categorical ground-truth answers for responses Xiong et al. (2024); Li et al. (2024); Wu et al. (2024); Nori et al. (2023), which fail to capture the complexities and risks associated with open-ended response generations. Where open-ended answers are evaluated, n-gram-based metrics such as BLEU Papineni et al. (2002) or ROUGE Lin (2004), historically used for machine translation, have been used Chen et al. (2019). However, these evaluations have been criticized for failing to capture the nuanced requirements of clinical QA, and even transformer-based metrics such as BertScore Zhang et al. (2019) have faced numerous semantic limitations Dada et al. (2024).
A key feature of these risks in the context of open-ended clinical QA is their non-binary nature (i.e., an answer is not simply "safe" or "unsafe" on a single axis). Consequently, the gold standard for assessing clinical inappropriateness remains human evaluation. For instance, Google’s work in clinical QA involved both clinicians and lay individuals, labeling responses based on various axes such as the likelihood and severity of harm, alignment with scientific consensus, and helpfulness Singhal et al. (2023a). Similarly, other studies have employed multi-axis evaluations with human clinicians to assess the overall appropriateness of responses for open-ended clinical QA Mukherjee et al. (2024); Singhal et al. (2023b); Zakka et al. (2024); Chowdhury et al. (2023).
@@ 2.2 Current RAG metrics
Evaluating RAG systems presents challenges due to their hybrid structure and the overall quality of the output often depends on multiple components within these systems. While attempts have been made to assess the overall quality of responses using deterministic methods Liu et al. (2023); Lyu et al. (2024), most current evaluation metrics for RAG systems use an ensemble of component-level assessments, the majority of which leverage LLMs as judges Yu et al. (2024). Broadly, RAG pipelines and the axes used to assess their performance can be broken down into the following components.
Retrieval component
The retrieval component is responsible for extracting relevant context from a knowledge source to match a given query.
  • Relevance (Context $\leftrightarrow$ Question): Measures how well the retrieved context matches the query’s information needs.
  • Accuracy (Relevant Context $\leftrightarrow$ Context Candidates): Assesses the accuracy of retrieved context compared to a set of candidates.
Generation component
where a model generates a response using information from the given context
  • Relevance (Response $\leftrightarrow$ Question): Evaluates the alignment of generated responses with the question’s intent.
  • Faithfulness / Groundedness (Response $\leftrightarrow$ Context): Assesses whether generated responses accurately reflect the retrieved context.
  • Correctness (Response $\leftrightarrow$ Sample Response): Measures the factual correctness of generated responses against a sample or standard response.
These component evaluations have been variably implemented with popular tools including TruEra’s RAG Triad Trulens (2023), and LangChain Bench LangChain . Additionally, LLM-as-a-judge-based frameworks like RAGAS Es et al. (2023), and ARES Saad-Falcon et al. (2023) have popularized common evaluation triads to capture possible permutations of the above components. Refer to Appendix A for an example of how the three components of the RAG system can be judged by LLMs, using RAGAS as an example.
@@ 2.2.1 Limitations of current metrics
The established methods to measure Faithfulness break down a model’s response into granular statements and then evaluate each statement’s consistency with the context Es et al. (2023). This approach aims to create more focused assertions that consider the context of both the question and the answer. It is particularly advantageous when answers are short and lack sufficient context when reviewed in isolation, as demonstrated in Figure 7. However, in the context of clinical conversations, this approach has the following shortcomings:
  • Summarizing the response into statements often neglects the clinical nuances in the original response (Figure 8).
  • Creating statements from both the patient’s question and the agent’s answer prevents the independent review of the agent’s answer concerning the context. This is especially problematic when the combined statement contains factually incorrect information (Figure 9).
  • Dialogue agents, particularly in clinical settings, are prompted to respond empathetically and conversationally. Statements constructed from the agent’s acknowledgments and questions, such as those meant to clarify or follow up on the patient’s queries or concerns, are penalized by existing faithfulness definitions (Figure 10).
@@
Evaluating answer relevance is critical in QA systems to ensure generated responses align with query’s intent. However, most current definitions focus on lexical or semantic similarity between the question and the response [Siriwardhana et al. (2023); Es et al. (2023)]. This has a number of drawbacks:
  • It overemphasizes surface-level topic matching without accounting for deeper contextual understanding.
  • It fails to account for whether a context is appropriate given a clinical context.
  • It does not handle "non-answers", meaning it struggles to determine when a system correctly discerns that a question falls outside its scope of relevance or when there is insufficient information to provide a safe and accurate response.
Furthermore, these metrics often do not reflect when the system appropriately refuses to address the question. This is critical in a clinical setting as clinicians, and similarly clinical QA systems must stay within their scope of practice.
@@ 3.1 Deriving metrics towards a safety case
In order to align our framework towards the evidence required to demonstrate if a clinical system is safe, we sought inspiration from published safety engineering frameworks - namely the Safety Assurance of autonomous systems in Complex Environments (SACE) guidance Hawkins et al. (2022). Structured safety engineering approaches have been applied towards the assurance of high-integrity autonomous systems (AS) such as maritime vessels Nakashima et al. (2023), automotive Rahman et al. (2023); Hunter et al. (2024), aerospace Torens et al. (2022), and healthcare domains Jia et al. (2022); Festor et al. (2022).
The SACE framework, in particular, provides a process to systematically integrate safety assurance into the development of AS whilst considering the system and its environment. Whilst we do not report all artifacts from the process in its entirety, we highlight a few key steps in this process that have been applied towards ASTRID’s design. Specifically, we considered the principles of:
  • Operating Context Assurance: Identifying the different clinical scenarios a patient might pose to a QA agent (see Figure 2).
  • Hazardous Scenario Identification: Analyzing how RAG systems could behave in hazardous ways within these scenarios.
  • Safe Operating Concept Assurance: Defining how an ideal system should behave in response to different queries.
  • Out-of-Context Operation Assurance: Determining how the system should respond when a question falls outside the scope.
Figure 2: Clinical Operating Contexts that face a clinical QA agent.
The clinical context is essential for determining appropriateness. For instance, the question, "Is it normal to have stomach cramps and vomiting?" would be irrelevant during a follow-up appointment for routine eye surgery, and the system should not to respond. However, if posed by a patient who has just been discharged after bowel surgery, it is not only highly relevant but also critical that the system provides a response. This is illustrated in Figure 3.
These concepts were outlined in a workshop where the dataset of real-world questions posed by patients to a voice-based AI conversational question were reviewed. The workshop consisted of AI developers, a clinician, and a safety practitioner (summarized in Figure 14). The analysis served as a bridge between subjective clinical assessments of harm and helpfulness and component-level validation scenarios for appropriate system performance. From the subsequent hazard analysis and the definition of safe operating requirements, it became clear that existing RAG evaluation metrics do not adequately capture key clinical risks in a conversational QA setting.
@@ 3.2 A novel set of metrics and a framework to assess safety risks
Current RAG QA metrics do not correlate to clinical risks, and have varying levels of validation against human evaluations, often performing poorly in conversational contexts. To our knowledge, no efforts have been made to connect QA system performance—measured by these metrics—with real-world clinician grading of clinical harms. For developers to meaningfully assess whether a clinical RAG QA system meets safe operating concepts, a framework is needed that is validated for clinical use, scalable, and capable of accounting for nuanced clinical contexts.
We propose a novel Automated and Scalable TRIaD (ASTRID) analysis framework for RAG-based clinical QA systems. ASTRID consists of three reference-free LLM-based metrics: Refusal Accuracy (RA), Conversational Faithfulness (CF) and Context Relevance (CR) (Figure 4). In the subsequent sections, we will illustrate how we validate each of the metrics and the overall framework based on a real world data from patients speaking to clinical conversational agents, augmented to ensure sufficient test case coverage.
Figure 4: ASTRID - an Automated and Scalable TRIaD for evaluating clinical QA systems leveraging RAG - consisting of three metrics: Context Relevance, Refusal Accuracy, and Conversational Faithfulness assessed within a clinical context.
@@ 3.2.1 Conversational Faithfulness (CF)
Evaluating the alignment of a response with the provided information is crucial for QA systems using RAG. Existing metrics fail to capture the complexities of conversational agents in clinical settings. We propose a newly-defined metric, Conversational Faithfulness, to address this gap.
Given an answer-context pair, CF is defined as the proportion of information-containing sentences that are faithful to the context. To calculate CF, we employ the following steps:
@@
    We categorize sentences in the response as either "information-containing" or "not information-containing" and extract the information-containing sentences.
@@
    We determine whether the information-containing sentences are grounded in context.
The prompts used to execute these steps are provided in Appendix 11. Finally, CF is calculated as follows:
$\texttt{CF}=\begin{cases}1,&\text{if }N=0\\ \frac{Y}{N},&\text{if }R\geq 0\\ 0,&\text{otherwise}\end{cases}$ (1)
where:
Y = Number of information-containing statements grounded in context
N = Total number of information-containing statements
R = N - (Y + Number of information-containing statements not grounded in context")
@@ 3.2.2 Refusal Accuracy (RA)
As discussed in previous sections, ensuring that a QA system appropriately declines to respond when a question is unanswerable or contextually inappropriate is critical in clinical settings, particularly for LLM-powered systems prone to generating ungrounded responses. Existing metrics do not capture this behavior, prompting us to introduce Refusal Accuracy.
Refusal Accuracy measures a system’s ability to withhold a response when no relevant information is available. It is evaluated using binary labels indicating whether the system appropriately refuses to answer. The prompt used for this assessment is provided in Appendix 13.